# From Planetesimals to Dwarf Planets by Pebble Accretion

### v1

@@ 5.4 The definition of a dwarf planet
The International Astronomical Union (IAU) defines a dwarf planet as “an object in orbit around the Sun that is large enough to pull itself into a nearly round shape but has not been able to clear its orbit of debris”  [^12] https://www.iau.org/public/themes/pluto/. While planetesimals formed through SI can exhibit irregular shapes (as exemplified by Arrokoth’s bi-lobed structure), significant pebble accretion would naturally lead to more spherical bodies through the slow settling of material and the isotropy of the accretion process. Therefore, a body grown by pebble accretion with $m_{\mathrm{final}}\gg m_{\mathrm{init}}$ is expected to have a nearly round shape. Specifically, if we set the mass gain factor at 10, $m_{\text{final}}=10m_{\text{init}}$, our simulations indicate that TNOs with masses greater than ${\sim}10^{-4}m_{\oplus}$ (or $6\times 10^{23}$ g, comparable to Orcus) would fulfill the IAU requirement of dwarf planet (Figure 7). Future stellar occultations of TNOs are essential to test this hypothesis.
@@ 6 Conclusions
In this study, we investigated a scenario in which high-mass trans-Neptunian objects (TNOs) accreted pebbles before being implanted into their current locations. The characteristics of the TNO size distribution suggest that the dynamically hot bodies featured processing at the high-mass end, where the distribution significantly flattens (see Figure 2). This finding aligns with the pebble accretion mechanism, which operates preferentially on the most massive bodies—those exceeding the critical mass threshold $m_{\ast}$ (Figure 4). Focusing on a carefully selected sample of dynamically hot TNOs within 48 au, for which we argue the population is complete ($H_{r}<5$ for bodies in the Minor Planet Survey and $H_{r}<8.3$ for bodies in OSSOS++), we performed an MCMC analysis to best reproduce the observed present-day size distribution. This analysis allowed us to constrain the key parameters that describe the pebble accretion process, specifically the pebbles aerodynamic size ($\tau_{s}$) and the pebble-planetesimal velocity ($\Delta v$).
@@
    The TNO size distribution tightly constrains the combination of the TNO-pebble relative velocity $\Delta v$ and pebble aerodynamic size $\tau_{s}$ in the primordial belt (Table 2 and Equation (15)). This corresponding threshold mass $m_{\ast}$, implies that bodies above ${\sim}10^{22}\,\mathrm{g}$ have enjoyed significant growth by pebble accretion (see Figure 7). While simulations featuring $\tau_{s}$ smaller than ${\sim}10^{-3}$ would formally fit the data, they imply very long accretion times, which rival the lifetime of the disk. Bodies exceeding $m_{\ast}$ increase their mass by many tenfolds, becoming Dwarf Planets (see Figure 7).
@@
    The data can be fitted with either a laminar or a turbulent velocity field. There is at presence no evidence for a second pebble size. This stems mainly from the lack of bodies in the high-mass tail of the TNO distribution, which render the fit insensitive to this part.
@@
    TNOs likely did not form in a laminar disk. In such an environment, the radial pressure gradient would imply the particle aerodynamic size $\tau_{s}\lesssim 10^{-3}$. Such small pebbles couple too well to the gas for pebble accretion to be effective. It would have made the process too slow in comparison to disk lifetimes. In addition, pebble accretion in smooth disks would put enormous demands on the total mass of pebbles that must have passed through the primordial belt (Equation (21); most of which do not get accreted). These concerns would be mitigated, however, in a cold disk environment.
@@
    Most likely, pebble accretion proceeded locally, under conditions where pebbles are entrained by pressure and could not escape from the birth ring. In this case, $\eta=0$ and the relative velocity is due to turbulence. These conditions would resemble the observed continuum mm-emission from continuum rings by ALMA. The inferred ($\sigma,\tau_{s}$) relation constrained from this work can be combined with the constraints the extent of the rings provide to give a preferred pebble aerodynamic size of $\tau_{s}\approx 10^{-2}$. Pebble accretion could have been complete in ${\sim}0.1{-}1$ Myr.
Future surveys like CLASSY (Fraser et al., 2022) and especially the advent of the Vera C. Rubin Observatory (LSST, Collaboration et al. 2021) will greatly expand the TNO inventory by ten folds. This will allow us to further constrain the pebble accretion process and deepen our understanding of planet formation in greater detail.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China under grant No. 12250610189, 12233004, and 12473065. Y.H. acknowledges funding support from Tsinghua University and National Observatory of Japan. The authors appreciate the referee for a thoughtful report, B. Gladman and E. Kokubo for useful discussions, and J.-M. Petit for providing the original OSSOS++ data.
@@ Appendix A TNO masses
with $\nu$ the albedo. For the cold belt, we take $\nu=0.15$ and $\rho_{\bullet}=1\,\mathrm{g\,cm^{-3}}$ (Petit et al. 2023), and with $\beta=0.252$ and $H_{B}=8.1$ (Kavelaars et al. 2021), the breaking mass amounts to $m_{\mathrm{bkr}}=2.3\times 10^{20}\,\mathrm{g}$ and the prefactor ($N_{\mathrm{pre}}$) for the cold belt $N_{\mathrm{cold}}=1.9\times 10^{4}$. For reference, the total mass, integrating Equation (2), amounts to
@@
which equals $2.0\times 10^{-3}\,m_{\oplus}$ for the cold belt. Also, the characteristic mass of the population, defined as the ratio of the second-to-first moment of the mass distribution, is $m_{\ast}=2\Gamma[(2-\alpha)/\gamma]/\Gamma[(1-\alpha)/\gamma]=3.97m_{\mathrm{bkr}}$.
For the hot belt, Petit et al. (2023) have shown that the shape of the magnitude distribution is the same as the cold belt with the number elevated by a factor 2.2. This would nevertheless amount to a different shape for the mass distribution (Equation (2)) since the albedos of the hot and cold belts differ, as is commonly assumed. Specifically, we take $\nu=0.08$ and $\rho_{\bullet}=1\,\mathrm{g\,cm^{-3}}$ for bodies in P48. However, at high mass it is known that the mass-magnitude relationship represented by Equations (24) and (28) breaks (see Figure 9). Several TNOs have mass estimates (those in P48 are listed in Table 4) and we can use these to fit a mass-radius relationship for the high mass end:

### v2

@@ 5.4 The definition of a dwarf planet
The International Astronomical Union (IAU) defines a dwarf planet as “an object in orbit around the Sun that is large enough to pull itself into a nearly round shape but has not been able to clear its orbit of debris.”  [^12] https://www.iau.org/public/themes/pluto/. While planetesimals that formed through SI can exhibit irregular shapes (as exemplified by Arrokoth’s bi-lobed structure), significant pebble accretion would naturally lead to more spherical bodies through the slow settling of material and the isotropy of the accretion process. Therefore, a body grown by pebble accretion with $m_{\mathrm{final}}\gg m_{\mathrm{init}}$ is expected to have a nearly round shape. Specifically, if we set the mass gain factor at 10, $m_{\text{final}}=10m_{\text{init}}$, our simulations indicate that TNOs with masses greater than ${\sim}10^{-4}m_{\oplus}$ (or $6\times 10^{23}$ g, comparable to Orcus) would fulfill the IAU requirement of dwarf planet (Figure 7). Future stellar occultations of TNOs are essential to test this hypothesis.
@@ 6 Conclusions
In this study, we investigated a scenario in which high-mass trans-Neptunian objects (TNOs) accreted pebbles before being implanted into their current locations. The characteristics of the TNO size distribution suggest that the dynamically hot bodies featured processing at the high-mass end, where the distribution significantly flattens (see Figure 2). This finding aligns with the pebble accretion mechanism, which operates preferentially on the most massive bodies, those exceeding the critical mass threshold $m_{\ast}$ (Figure 4). Focusing on a carefully selected sample of dynamically hot TNOs within 48 au, for which we argue the population is complete ($H_{r}<5$ for bodies in the Minor Planet Survey and $H_{r}<8.3$ for bodies in OSSOS++), we performed an MCMC analysis to best reproduce the observed present-day size distribution. This analysis allowed us to constrain the key parameters that describe the pebble accretion process, specifically the pebble aerodynamic size ($\tau_{s}$) and the pebble-planetesimal velocity ($\Delta v$).
@@
    The TNO size distribution tightly constrains the combination of the TNO-pebble relative velocity $\Delta v$ and pebble aerodynamic size $\tau_{s}$ in the primordial belt (Table 2 and Equation (15)). This corresponding threshold mass $m_{\ast}$, implies that bodies above ${\sim}10^{22}\,\mathrm{g}$ have enjoyed significant growth by pebble accretion (see Figure 7). While simulations featuring $\tau_{s}$ smaller than ${\sim}10^{-3}$ would formally fit the data, they imply very long accretion times, which rival the lifetime of the disk. Bodies exceeding $m_{\ast}$ increase their mass several tenfold and become dwarf planets (see Figure 7).
@@
    The data can be fitted with either a laminar or a turbulent velocity field. There is at present no evidence of a second pebble size. This stems mainly from the lack of bodies in the high-mass tail of the TNO distribution, which render the fit insensitive to this part.
@@
    TNOs likely did not form in a laminar disk. In such an environment, the radial pressure gradient would imply the particle aerodynamic size $\tau_{s}\lesssim 10^{-3}$. Such small pebbles couple too well to the gas for pebble accretion to be effective. It would have made the process too slow in comparison to disk lifetimes. In addition, pebble accretion in smooth disks would put enormous demands on the total mass of pebbles that must have passed through the primordial belt (Equation (21), most of which do not get accreted). These issues would be mitigated, however, in a cold disk environment.
@@
    Most likely, pebble accretion proceeded locally, under conditions where pebbles are entrained by pressure and could not escape from the birth ring. In this case, $\eta=0$ and the relative velocity is due to turbulence. These conditions would resemble the observed continuum millimeter emission from continuum rings by ALMA. The inferred ($\sigma,\tau_{s}$) relation from this work can be combined with the constraints that the extent of the rings provide to give a preferred pebble aerodynamic size of $\tau_{s}\approx 10^{-2}$. Pebble accretion could have been complete in ${\sim}0.1{-}1$ Myr.
Future surveys like CLASSY (Fraser et al., 2022) and especially the advent of the Vera C. Rubin Observatory (LSST, Collaboration et al. 2021) will greatly expand the TNO inventory by many times over. This will allow us to further constrain the pebble accretion process and deepen our understanding of planet formation in greater detail.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China under grant No. 12473065, 12233004, and 12250610189. Y.H. acknowledges funding support from Tsinghua University and National Observatory of Japan. The authors appreciate the referee for a thoughtful report, B. Gladman and E. Kokubo for useful discussions, and J.-M. Petit for providing the original OSSOS++ data.
@@ Appendix A TNO masses
with $\nu$ the albedo. For the cold belt, we take $\nu=0.15$ and $\rho_{\bullet}=1\,\mathrm{g\,cm^{-3}}$ (Petit et al. 2023), which, with $\beta=0.252$ and $H_{B}=8.1$ (Kavelaars et al. 2021), amounts to a breaking mass of $m_{\mathrm{bkr}}=2.3\times 10^{20}\,\mathrm{g}$ and a prefactor for the cold belt $N_{\mathrm{cold}}=N_{\mathrm{pre}}=1.9\times 10^{4}$. For reference, the total mass, integrating Equation (2), amounts to
@@
which equals $2.0\times 10^{-3}\,m_{\oplus}$ for the cold belt. In addition, the characteristic mass of the population, defined as the ratio of the second to first moment of the mass distribution, is $m_{\ast}=2\Gamma[(2-\alpha)/\gamma]/\Gamma[(1-\alpha)/\gamma]=3.97m_{\mathrm{bkr}}$.
For the hot belt, Petit et al. (2023) have shown that the shape of the magnitude distribution is the same as the cold belt with the number elevated by a factor of 2.2. This would nevertheless amount to a different shape for the mass distribution (Equation (2)) since the albedos of the hot and cold belts differ, as is commonly assumed. Specifically, we take $\nu=0.08$ and $\rho_{\bullet}=1\,\mathrm{g\,cm^{-3}}$ for bodies in P48. However, at high mass it is known that the mass-magnitude relationship represented by Equations (24) and (28) breaks (see Figure 9). Several TNOs have mass estimates (those in P48 are listed in Table 4) and we can use them to fit a mass–radius relationship for the high-mass end: